# Energy Extraction From the Kerr-Bertotti-Robinson Black Hole via Magnetic Reconnection under Circular Plasma and Plunging Plasma

### v1

@@ I Introdution
Black holes are the most fascinating objects in the universe. Since the detection of gravitational waves by Laser Interferometer Gravitational-wave Observatory (LIGO) 41 and the imaging of the black hole shadow by the Event Horizon Telescope (EHT) 42 , physicists have gained a deeper understanding of black holes. Previous research has demonstrated that matter surrounding a black hole significantly influences the trajectories of particles and the appearance of its shadow image. This is particularly evident in the EHT’s published analyses of polarized imaging around the supermassive black hole M87* 43 and the polarized imaging of the Galactic center black hole Sagittarius A* 44 , which confirmed the presence of magnetic fields around these black holes. Therefore, studying the properties of magnetized black holes is important both theoretically and astronomically. Currently, numerous studies have explored the physical properties of black holes immersed in magnetic fields. Most of these build upon the Schwarzschild-Melvin black hole solution, which has subsequently been generalized to include rotation, electric charge, and other parameters 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 ; 21 ; 22 ; 23 ; 7 ; 8 ; 9 ; 11 . However, these spacetimes have some drawbacks. For instance, the magnetic field decreases far away from the black hole, the geodesics cannot escape to infinity 1 , and the algebraic type of the spacetime is $I$ 23 . Recently, a new class of exact black hole solutions immersed in a magnetic field has been discovered 1 . This black hole is referred to as the Kerr-Bertotti-Robinson black hole. Crucially, test particles can, in principle, escape to spatial infinity within this spacetime, and its algebraic type is $D$ 1 . Many properties of this new black hole remain unclear. This paper aims to explore how energy can be extracted from this black hole via magnetic reconnection in the circular orbit region and plunging region.
@@
Extracting energy from black holes has long been a topic of great interest to physicists. Penrose discovered in 1969 that energy could be extracted from rotating black holes, a process now known as the Penrose process 5 . Later, Wald pointed out 24 that realizing the Penrose process in practice is extremely difficult. Building on Penrose’s pioneering work, many physicists have discovered other mechanisms for energy extraction, such as superradiant scattering 25 , the Blandford-Znajek process 26 , among others. Comisso and Asenjo also discovered a highly efficient mechanism for energy extraction by employing the method of magnetic reconnection 3 (for related earlier work, see 45 ; 46 ). Furthermore, in some cases, its power exceeds that of the Blandford-Znajek process. Consequently, this work has been rapidly extended to other rotating black holes 2 ; 6 ; 27 ; 28 ; 29 ; 30 ; 31 ; 32 ; 33 ; 34 ; 35 ; 36 ; 37 . Currently, research on the magnetic reconnection mechanism primarily focuses on the circular orbit region. Recently, reference 13 proposed studying magnetic reconnection and energy extraction in the plunging region. Specifically, plasma initially performs circular motion outside the innermost stable circular orbit (ISCO). However, since circular orbits within the ISCO are unstable, the plasma plunges inward starting from the ISCO. References 13 ; 38 ; 39 ; 40 have already investigated magnetic reconnection and energy extraction in the plunging region. The results demonstrate that energy extraction from the black hole is indeed possible in the plunging region. More importantly, both the power and efficiency of energy extraction in the plunging region are significantly higher than those in the circular orbit region.
@@ II Introduction to Kerr-Bertotti-Robinson Spacetime
In Boyer-Lindquist (BL) coordinates with geometric units ($c=G=1$), the metric expression for the Kerr-Bertotti-Robinson spacetime is given by 1
@@
We consider the equations of motion for particles confined to the equatorial plane ($\theta=\pi/2$). The normalization condition for the geodesic equations is given by
@@
$R(r)=\Omega^{4}\left(\left[E(r^{2}+a^{2})-aL\right]^{2}-\Omega^{2}(aE-L)^{2}\Delta-\varepsilon\Delta r^{2}\right).$ (18)
@@
For energy extraction, only prograde orbits need to be considered. The Keplerian angular velocity for particles in prograde orbits is given by 2
@@ III.1 Magnetic Reconnection Process in Circular Orbits
The magnetic reconnection process in circular orbits follows the Comisso-Asenjo process 3 . We first review the Comisso-Asenjo process. Observations are conducted using the Zero Angular Momentum Observer (ZAMO) reference frame 4 . In this frame, the spacetime interval is expressed as
@@
For a relativistically hot plasma characterized by the equation of state $w=4p$, energy extraction, which is analogous to the Penrose mechanism 5 , requires the simultaneous satisfaction of two criteria
@@
We plot the variation of $e_{+}^{\infty}$ and $e_{-}^{\infty}$ with azimuthal angle $\xi$ for different values of $B$ in Fig. 3. Here $r$ denotes the dominant reconnection radical location. Without loss of generality, we set $a=0.99$, $w=0.001$, and $r=1.5$. Solid lines represent $e_{+}^{\infty}$, dashed lines represent $e_{-}^{\infty}$, with colors blue, red, green, orange, magenta, and brown corresponding to $B=0.05,0.1,0.2,0.3,0.4,0.5$ respectively. The right panel shows a zoomed-in view of the left panel near $e_{-}^{\infty}=0$.
(a) $B$=0.05
(b) $B$=0.1
(c) $B$=0.2
(d) $B$=0.3
(e) $B$=0.4
(f) $B$=0.5
@@ III.2 Power and Efficiency of Energy Extraction in Circular Orbits
For collisionless conditions, $U_{in}\approx 0.1$, for collisional conditions, $U_{in}\approx 0.01$. This work considers the collisionless case. $A_{in}$ is the cross-sectional area of the inflowing plasma can be expressed as
@@
From Table 1, it can be observed that the power of Kerr-Bertotti-Robinson black holes consistently falls below that of Kerr black holes. This discrepancy widens with increasing magnetic field strength. The difference from the Kerr black hole is somewhat smaller when $a=0.99$. An anomalous behavior emerges at $a=0.99$ for $B=0.45$ and 0.50.50.5, where the power suddenly increases. This reversal stems from proximity to extremal black holes, which facilitates energy extraction, however still fails to surpass Kerr black hole performance. This indicates that magnetic fields hinder energy extraction, with stronger fields imposing greater suppression, consistent with findings in reference 6 . Notably, while magnetic fields still impede power output, the novel Kerr-Bertotti-Robinson solution enables more efficient energy extraction than Kerr-Melvin black holes. As $B$ increases, Kerr-Bertotti-Robinson’s power advantage over Kerr-Melvin grows, demonstrating its superior energy-extraction capability in strong magnetic fields. Compared to the $a=0.99$ case, Kerr-Bertotti-Robinson black holes at $a=0.97$ exhibit even greater power superiority over Kerr-Melvin solutions. This suggests that away from extremal limits, Kerr-Bertotti-Robinson black holes offer enhanced advantages for energy extraction.
@@ IV.1 Magnetic Reconnection Process in the Plunging Region
Next, we consider the scenario in the plunging region, as illustrated in Fig. 7.
Figure 7: Schematic diagram of circular orbits and the plunging region.
Plasma initially follows a circular trajectory just outside ISCO. However, circular orbits become unstable within the ISCO radius. Consequently, any disturbance triggers radial motion, causing the plasma to plunge inwards starting precisely at the ISCO. This inner domain is termed the plunging region. Here, the presence of significant radial velocity invalidates the Keplerian velocity formula given in (27), as that expression accounts solely for the azimuthal component. For comprehensive computational details regarding the plunging region, refer to reference 13 . Here, we briefly review the energy extraction process in the plunging region. We continue to utilize the convenient ZAMO frame. The transformation of four-velocity components between the Boyer-Lindquist (BL) coordinates and the ZAMO frame is expressed as
@@
[…]
@@ IV.2 Power and Efficiency of Energy Extraction in the Plunging Region
Within the plunging region, we similarly compare the power outputs among Kerr-Bertotti-Robinson, Kerr, and Kerr-Melvin black holes in Table 2. First, comparing Kerr and Kerr-Bertotti-Robinson black holes. Table 2 reveals that magnetic fields still impede energy extraction, with stronger fields causing greater suppression. However, the power ratios here are more favorable than in circular orbits, except for $a=0.99$ at $B=0.5$. The anomalous rise at $a=0.99$ for $B=0.45$ and 0.50.50.5 stems from extremal black hole effects, with both plunging and circular orbit regions exhibiting increased power near $B=0.45$. This reaffirms that proximity to extremal black holes facilitates energy extraction. Next, comparing Kerr-Bertotti-Robinson and Kerr-Melvin black holes. The variation pattern in the plunging region resembles that in circular orbits. The novel Kerr-Bertotti-Robinson solution consistently outperforms Kerr-Melvin black holes, demonstrating greater advantages in strong magnetic fields and particularly when not near extremal limits.
@@
$P_{\mathrm{K}}/P_{\mathrm{Kerr}}$
@@ V Conclusion
Recent years, more and more astronomical observations indicates the presence of magnetic fields around black holes. Black hole solutions incorporating magnetic fields are therefore crucial for understanding significant astrophysical phenomena. Existing descriptions of magnetized black holes, such as Melvin-like black holes, has many drawbacks. For example, magnetic fields diminish at large distances, geodesics in these spacetimes cannot escape to infinity, and their algebraic type is $I$, rendering them unrealistic models for astrophysical magnetized black holes. The exact solution recently proposed in Ref. 1 circumvents these issues. Many properties of this black hole remain unexplored. So we investigated energy extraction from the Kerr-Bertotti-Robinson black hole via the magnetic reconnection mechanism and compared the results with those of the Kerr-Melvin black hole. Specifically, we conducted studies in both the circular orbits region and the plunging region and compared the corresponding results.

### v2

@@ I Introdution
Black holes are one of the most mysterious celestial objects in the universe. Since the detection of gravitational waves by Laser Interferometer Gravitational-wave Observatory (LIGO) 41 and the imaging of the black hole shadow by the Event Horizon Telescope (EHT) 42 , physicists have gained a deeper understanding of black holes. Previous research has demonstrated that matter surrounding a black hole significantly influences the trajectories of particles and the appearance of its shadow image. This is particularly evident in the EHT’s published analyses of polarized imaging around the supermassive black hole M87* 43 and the polarized imaging of the Galactic center black hole Sagittarius A* 44 , which confirmed the presence of magnetic fields around these black holes. Therefore, studying the properties of magnetized black holes is important both theoretically and astronomically. Currently, numerous studies have explored the physical properties of black holes immersed in magnetic fields. Most of these build upon the Schwarzschild-Melvin black hole solution, which has subsequently been generalized to include rotation, electric charge, and other parameters 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 ; 21 ; 22 ; 23 ; 7 ; 8 ; 9 ; 11 ; Barrientos:2024pkt . However, these spacetimes have some drawbacks. For instance, the magnetic field decreases far away from the black hole, the geodesics cannot escape to infinity 1 , and the algebraic type of the spacetime is $I$ 23 . Recently, a new class of exact black hole solutions immersed in a magnetic field has been discovered 1 . This black hole is referred to as the Kerr-Bertotti-Robinson black hole. Crucially, test particles can, in principle, escape to spatial infinity within this spacetime, and its algebraic type is $D$ 1 . Type $D$ spacetimes often possess additional symmetries (Killing vector), which typically allow for the separation of variables in null geodesic equations. This significantly simplifies both the precise analytical and numerical studies of particle orbits. In contrast, type I spacetimes generally lack such symmetries, resulting in the inability to separate variables in the geodesic equations. Many properties of this new black hole remain unclear. This paper aims to explore how energy can be extracted from this black hole via magnetic reconnection in the circular orbit region and plunging region.
@@
Extracting energy from black holes has long been a topic of great interest to physicists. Penrose discovered in 1969 that energy could be extracted from rotating black holes, a process now known as the Penrose process 5 . A particle of energy $E$ falls freely from infinity into a black hole’s ergosphere and splits. From an observer at infinity’s viewpoint, a negative-energy fragment enters the event horizon, allowing a positive-energy fragment to escape along a geodesic. This constitutes energy extraction. Later, Wald pointed out 24 that realizing the Penrose process in practice is extremely difficult. Building on Penrose’s pioneering work, many physicists have discovered other mechanisms for energy extraction, such as superradiant scattering 25 , the Blandford-Znajek process 26 , among others. Comisso and Asenjo also discovered a highly efficient mechanism for energy extraction by employing the method of magnetic reconnection 3 (for related earlier work, see 45 ; 46 ). Furthermore, in some cases, its power exceeds that of the Blandford-Znajek process. Consequently, this work has been rapidly extended to other rotating black holes 2 ; 6 ; 27 ; 28 ; 29 ; 30 ; 31 ; 32 ; 33 ; 34 ; 35 ; 36 ; 37 ; Fan:2024rsa ; Fan:2024fcy . Currently, research on the magnetic reconnection mechanism primarily focuses on the circular orbit region. Recently, reference 13 proposed studying magnetic reconnection and energy extraction in the plunging region. Specifically, plasma initially performs circular motion outside the innermost stable circular orbit (ISCO). However, since circular orbits within the ISCO are unstable, the plasma plunges inward starting from the ISCO. References 13 ; 38 ; 39 ; 40 have already investigated magnetic reconnection and energy extraction in the plunging region. The results demonstrate that energy extraction from the black hole is indeed possible in the plunging region. More importantly, for moderate to low black hole spin values $a$ (where a significant plunging region exists inside the ISCO), both the power and efficiency of energy extraction in this plunging region are typically higher than those achievable from circular orbits in the equatorial plane.
@@ II Introduction to Kerr-Bertotti-Robinson Spacetime
In Boyer-Lindquist (BL) coordinates with geometric units ($c=G=1$), the metric expression for the Kerr-Bertotti-Robinson spacetime is given by 1 ; 52 ; 53
@@
We follow reference 3 , where a simplifying assumption is made that the magnetic reconnection process is modeled in the current sheet located on the equatorial plane. Therefore, we restrict the particle motion to the equatorial plane. While this approach may limit the generality of our conclusions, reference 54 shows that the geodesic equations of this spacetime cannot be solved via variable separation for massive particles. This significantly increases the complexity of analyzing magnetic reconnection under general conditions. So, we consider the equations of motion for particles confined to the equatorial plane ($\theta=\pi/2$). The normalization condition for the geodesic equations is given by
@@
$R(r)=\left[E(r^{2}+a^{2})-aL\right]^{2}-\Omega^{2}(aE-L)^{2}\Delta-\varepsilon\Delta r^{2}.$ (18)
@@
For retrograde orbits, the photon sphere radius lies outside the ergosphere, making energy extraction impossible under circular orbits. Regarding the plunging region, although particles can enter the ergosphere from the ISCO, the energy density per enthalpy at infinity for both accelerated and decelerated plasma remains greater than zero, which still prevents energy extraction 13 . Therefore, for energy extraction, only prograde orbits need to be considered. The Keplerian angular velocity for particles in prograde orbits is given by 2
@@ III.1 Magnetic Reconnection Process in Circular Orbits
The magnetic reconnection process follows the Comisso-Asenjo framework 3 for equatorial orbits. We first review the Comisso-Asenjo process. Observations are conducted using the Zero Angular Momentum Observer (ZAMO) reference frame 4 . In this frame, the spacetime interval is expressed as
@@
For a relativistically hot plasma characterized by the equation of state $w=4p$, energy extraction, which is analogous to the Penrose particle-splitting mechanism 5 , requires the simultaneous satisfaction of two criteria
@@
We plot the variation of $e_{+}^{\infty}$ and $e_{-}^{\infty}$ with azimuthal angle $\xi$ for different values of $B$ in Fig. 3. Here $r$ denotes the dominant reconnection radial location. As explained in 3 , the equatorial current sheet is unstable to the plasmoid instability 55 ; 56 , leading to fragmentation into multiple X-points. Among these, the dominant reconnection X-point is the one located at the intersection of the separatrices that encompass the global reconnection flow. This is the point that determines the reconnection dynamics and is the one referred to by the distance $r$. Without loss of generality, we set $a=0.99$, $w=0.001$, and $r=1.5$. Solid lines represent $e_{+}^{\infty}$, dashed lines represent $e_{-}^{\infty}$, with colors blue, red, green, orange, magenta, and brown corresponding to $B=0.05,0.1,0.2,0.3,0.4,0.5$ respectively. The right panel shows a zoomed-in view of the left panel near $e_{-}^{\infty}=0$.
(a) BB=0.05
(b) BB=0.1
(c) BB=0.2
(d) BB=0.3
(e) BB=0.4
(f) BB=0.5
(g) Legend
@@ III.2 Power and Efficiency of Energy Extraction in Circular Orbits
For collisionless conditions, $U_{in}\approx 0.1$ 57 , for collisional conditions, $U_{in}\approx 0.01$ 58 . This work considers the collisionless case. $A_{in}$ is the cross-sectional area of the inflowing plasma can be expressed as
@@
From Table 1, it can be observed that the power of Kerr-Bertotti-Robinson black holes consistently falls below that of Kerr black holes. This discrepancy widens with increasing magnetic field strength. The difference from the Kerr black hole is somewhat smaller when $a=0.99$. An anomalous behavior emerges at $a=0.99$ for $B=0.45$ and $0.5$, where the power suddenly increases. This reversal stems from proximity to extremal black holes, which facilitates energy extraction, however still fails to surpass Kerr black hole performance. This indicates that magnetic fields hinder energy extraction, with stronger fields imposing greater suppression, consistent with findings in reference 6 . Notably, while magnetic fields still impede power output, the novel Kerr-Bertotti-Robinson solution enables more efficient energy extraction than Kerr-Melvin black holes. As $B$ increases, Kerr-Bertotti-Robinson’s power advantage over Kerr-Melvin grows, demonstrating its superior energy-extraction capability in strong magnetic fields. Compared to the $a=0.99$ case, Kerr-Bertotti-Robinson black holes at $a=0.97$ exhibit even greater power superiority over Kerr-Melvin solutions. This suggests that away from extremal limits, Kerr-Bertotti-Robinson black holes offer enhanced advantages for energy extraction.
@@ IV.1 Magnetic Reconnection Process in the Plunging Region
Next, we consider the scenario in the plunging region. Plasma initially follows a circular trajectory just outside ISCO. However, circular orbits become unstable within the ISCO radius. Consequently, any disturbance triggers radial motion, causing the plasma to plunge inwards starting precisely at the ISCO. This inner domain is termed the plunging region. Here, the presence of significant radial velocity invalidates the Keplerian velocity formula given in (27), as that expression accounts solely for the azimuthal component. For comprehensive computational details regarding the plunging region, refer to reference 13 . Here, we briefly review the energy extraction process in the plunging region. We continue to utilize the convenient ZAMO frame. The transformation of four-velocity components between the Boyer-Lindquist (BL) coordinates and the ZAMO frame is expressed as
@@
$\begin{aligned} e_{\pm}^{\infty}&=\alpha\hat{\gamma}_{s}\gamma_{{out}}\left[\left(1+\beta^{\varphi}\hat{v}_{s}^{(\varphi)}\right)\pm v_{{out}}\left(\hat{v}_{s}+\beta^{\varphi}\frac{\hat{v}_{s}^{(\varphi)}}{\hat{v}_{s}}\right)\cos\xi\mp v_{{out}}\beta^{\varphi}\frac{\hat{v}_{s}^{(r)}}{\hat{\gamma}_{s}\hat{v}_{s}}\sin\xi\right]\\ &-\alpha\left(4\hat{\gamma}_{s}\gamma_{{out}}\left(1\pm\hat{v}_{s}v_{{out}}\cos\xi\right)\right)^{-1},\end{aligned}$ (38)
@@ IV.2 Power and Efficiency of Energy Extraction in the Plunging Region
Within the plunging region, we similarly compare the power outputs among Kerr-Bertotti-Robinson, Kerr, and Kerr-Melvin black holes in Table 2. First, comparing Kerr and Kerr-Bertotti-Robinson black holes. Table 2 reveals that magnetic fields still impede energy extraction, with stronger fields causing greater suppression. However, the power ratios here are more favorable than in circular orbits, except for $a=0.99$ at $B=0.5$. The anomalous rise at $a=0.99$ for $B=0.45$ and $0.5$ stems from extremal black hole effects, with both plunging and circular orbit regions exhibiting increased power near $B=0.45$. This reaffirms that proximity to extremal black holes facilitates energy extraction. Next, comparing Kerr-Bertotti-Robinson and Kerr-Melvin black holes. The variation pattern in the plunging region resembles that in circular orbits. The novel Kerr-Bertotti-Robinson solution consistently outperforms Kerr-Melvin black holes, demonstrating greater advantages in strong magnetic fields and particularly when not near extremal limits.
@@
$P_{\mathrm{}}/P_{\mathrm{Kerr}}$
@@ V Conclusion
Recent years, more and more astronomical observations indicates the presence of magnetic fields around black holes. Black hole solutions incorporating magnetic fields are therefore crucial for understanding significant astrophysical phenomena. Existing descriptions of magnetized black holes, such as Melvin-like black holes, has many drawbacks. For example, magnetic fields diminish at large distances, geodesics in these spacetimes cannot escape to infinity, and their algebraic type is $I$, rendering them unrealistic models for astrophysical magnetized black holes. The exact solution recently proposed in Ref. 1 circumvents these issues. Many properties of this black hole remain unexplored. So we investigated energy extraction from the Kerr-Bertotti-Robinson black hole via the Comisso-Asenjo process and compared the results with those of the Kerr-Melvin black hole. Specifically, we conducted studies in both the circular orbits region and the plunging region and compared the corresponding results.